OpenStudy (anonymous):

A solution containing 30% insecticide is to be mixed with a solution containing 50% insecticide to make 200 L of a solution containing 42% insecticide. How much of the 30% insecticide solution should be used?@ ßαbчღ@nyprincess2001

OpenStudy (anonymous):

let x L of 30% insecticide to be mixed with (200-x) L of 50% insecticide to make a 42% 200L insecticide solution so 0.3*x+0.5*(200-x)=0.42*200 100-(0.2*x)=84 0.2*x=16 x=80 L hence 80 L of 30% insecticide to be used.